# The legendary Mario and Luigi show us that whether you slap in the crossings as early as a warp pipe can shoot you or as late as the very last bend, the water system in Yoshi Hill comes out exactly the same!

### v1

@@ Abstract.
which are again counted by median Genocchi numbers! The graphs of (i) are known as terrain-like graphs and a similar bijection from (i) to (iii) was presented in [FR24] (but definitely more twisted than a Mario pipe maze!).” But he is suddenly interrupted in the middle of his wild musings…
@@
  3. (iii)
    permutations lower than $\tau$ in (strong) Bruhat order (recall that $\sigma\leq\tau$ if $\sigma$ admits a reduced expression which is a subword of a reduced expression of $\tau$).
@@
It is clear that any permutation $\sigma$ of (iii) satisfies that $\mu(i)\leq\sigma(i)\leq\nu(i)$ where $\mu(i)$ and $\nu(i)$ are the leftmost and rightmost customers reachable from the $i$th faucet (for example $\mu({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}19})={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11}$} and $\nu({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}19})={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}27}$ in Figure˜4). The reciprocal statement does not hold in general. Luckily, Toadsworth has hidden talents, and he already knew about this subtlety thanks to last year’s disasters on Lakitu Island, where $\tau=351624$, $\mu=111224$ and $\nu=355666$ (see Figure˜5 (a & b)). Back then, he had asked Mario and Luigi to handle several bill assignments. Our two plumbers managed to pull it off for the permutation $351624$ (see Figure˜5 (c)), but they eventually ended up with a full plate of pipe-spaghetti for the permutation $145236$ (see Figure˜5 (d & e)). Using that $\sigma\leq\tau$ in Bruhat order if and only if $\left\{k\leq i\;\middle|\;\sigma(k)\leq j\right\}\geq\#\left\{k\leq i\;\middle|\;\tau(k)\leq j\right\}$ for all $i,j\in[n]$, it is easy to check that the reciprocal statement does hold as soon as the top path has a unique peak (as in Yoshi hill) or the bottom path has a unique valley.
Figure 5. Lakitu Island. Here, we have $\tau=351624$, $\mu=111224$ and $\nu=355666$. The two permutations $\sigma=351624$ and $\sigma^{\prime}=145236$ both satisfies $\mu(i)\leq\sigma(i)\leq\nu(i)$ for all $i$, but only $\sigma$ is below $\tau$ in (strong) Bruhat order.
@@
I am also grateful to Federico Ardila and Christian Stump for discussions which happened at the CIRM research school “Beyond Permutahedra and associahedra” in December 2025.

### v2

@@ Abstract.
which are again counted by median Genocchi numbers!” But he is suddenly interrupted in the middle of his wild musings…
@@
  3. (c)
    permutations of $[m]$ lower than $\tau$ in (strong) Bruhat order (recall that $\sigma\leq\tau$ if $\sigma$ admits a reduced expression which is a subword of a reduced expression for $\tau$).
@@
Figure 5. Lakitu Island. Here, we have $\tau=351624$, $\mu=111224$ and $\nu=355666$. The two permutations $\sigma=351624$ and $\sigma^{\prime}=145236$ both satisfies $\mu(i)\leq\sigma(i)\leq\nu(i)$ for all $i\in[6]$, but only $\sigma$ is below $\tau$ in (strong) Bruhat order.
It is clear that any permutation $\sigma$ of (c) satisfies that $\mu(i)\leq\sigma(i)\leq\nu(i)$ for all $i\in[m]$, where $\mu(i)$ and $\nu(i)$ are the leftmost and rightmost customers reachable from the $i$th faucet (for example $\mu({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}19})={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11}$} and $\nu({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}19})={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}27}$ in Figure˜4).
The reciprocal statement does not hold in general. Luckily, Toadsworth has hidden talents, and he already knew about this subtlety thanks to last year’s disasters on Lakitu Island, where $\tau=351624$, $\mu=111224$ and $\nu=355666$ (see Figure˜5 (a & b)). Back then, he had asked Mario and Luigi to handle several bill assignments. Our two plumbers managed to pull it off for the permutation $\sigma=351624$ (see Figure˜5 (c)), but they eventually ended up with a full plate of pipe-spaghetti for the permutation $\sigma^{\prime}=145236$ (see Figure˜5 (d & e)). Both permutations $\sigma$ and $\sigma^{\prime}$ satisfies $\mu(i)\leq\sigma(i)\leq\nu(i)$ for all $i\in[6]$, but only $\sigma$ is below $\tau$ in (strong) Bruhat order.
Using that $\sigma\leq\tau$} in Bruhat order if and only if $\left\{k\leq i\;\middle|\;\sigma(k)\leq j\right\}\geq\#\left\{k\leq i\;\middle|\;\tau(k)\leq j\right\}$} for all $i,j\in[n]$}, it is easy to check that the reciprocal statement does hold as soon as the top path has a unique peak (as in Yoshi hill) or the bottom path has a unique valley.
@@
The crossing internally closed graphs (a) are also known as grounded rectangle graphs [JT19], hook graphs [Hix13], max point-tolerance graphs [CCF+17], $p$-box graphs [STC15], or non-jumping graphs [AFS19], while the crossing externally closed graphs (b) are also known as terrain-like graphs [FR21, FR24, AFS19]. A similar bijection from (b) to Dumont derangements (c) was presented in [FR24] (but definitely more twisted than a Mario pipe maze!), but I am not aware that the bijection from (a) to either (b) or (c) was observed earlier, even if the two classes of graphs (a) and (b) were compared in [AFS19].
@@
I am also grateful to Federico Ardila, Jean Cardinal, and Christian Stump for discussions which happened at the CIRM research school “Beyond Permutahedra and associahedra” in December 2025.